OpenStudy (anonymous):

anyone up to helping with my algebrA test this is the last chance i get to take it out the 3 and i failed the first 2 times trying

OpenStudy (unklerhaukus):

2/3=66.67%

OpenStudy (anonymous):

oh no that wasnt a problem it was a question i really need help doing my test i have taken it already twice i only have one chance to take it again

OpenStudy (unklerhaukus):

if you have algebra question i can help you with that

OpenStudy (anonymous):

okay thank u sooooo much

OpenStudy (anonymous):

y-23=-36

OpenStudy (anonymous):

is y 59?

OpenStudy (anonymous):

What is the goal of this question?

OpenStudy (anonymous):

to find what y is

OpenStudy (anonymous):

y-23=-36 i no if you subtract 23 from 59 you get -36

OpenStudy (anonymous):

right. You need to isolate y on one side of the equation. In this case, there is only one way to do that. You would add 23 to both sides.

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

you do that by doing the opposite of what is already shown. Since you have -23, you need to add 23. If it were positive 23, you would subtract 23

OpenStudy (unklerhaukus):

y-23 = -36 adding twenty three to both sides y-23 + 23 = -36 + 23

OpenStudy (anonymous):

okwhat about this one 8+x=-4 i no last night when i did 6+x=-2 i put in 8 and it turned out to be -8 so when i do 8+x=-4 im thinking the x is also -4?

OpenStudy (anonymous):

what do you need to do to the equation to isolate x on one side of it?

OpenStudy (anonymous):

yeah solve the equation

OpenStudy (anonymous):

i think its -4?

OpenStudy (anonymous):

why -4?

OpenStudy (unklerhaukus):

what did you get for y?

OpenStudy (anonymous):

not saying you're wrong about it being -4, I just want to see how you came to it.

OpenStudy (anonymous):

when i did one like this last night it was 6+x=-2 6-8=-2 x=-8 i missed this problem on the test because i didnt put the - sign

OpenStudy (anonymous):

8+x=-4 4-8=-4

OpenStudy (anonymous):

oooh wait is it 12? 8-12=-4

OpenStudy (anonymous):

If you have 8 + x = -4, you only need to subtract 8 from both sides

OpenStudy (anonymous):

x = - 4 - 8

OpenStudy (anonymous):

im confussed

OpenStudy (anonymous):

if you have 8 + x = - 4, what do you need to do to get x isolated on one side?